OpenStudy (anonymous):

What is the 7th term of the geometric sequence -1, 3, -9, ...?

OpenStudy (anonymous):

\[a _{1} = -1 \] This is the first term of the geometric sequence.

OpenStudy (anonymous):

r or the common ratio is \[r = 3/-1 = -3 \] \[r = -9/3 = -3 \] so \[r = -3 \]

OpenStudy (anonymous):

NOW USE THIS FORMULA \[a _{n} = a _{1} * r ^{n-1}\]

OpenStudy (anonymous):

where \[a _{1} = -1 \] and \[r = -3 \] and \[n = 7 \] since that is the number ur trying to find.

OpenStudy (anonymous):

Is it -729

OpenStudy (anonymous):

HAHAHA jesus lolol

OpenStudy (anonymous):

\[a _{7} = -1 * (-3)^{7-1}\] \[a _{7} = -1 * (-3)^{6}\] \[a _{7} = -1 * (729)\] \[a _{7} = -729\]

OpenStudy (anonymous):

I have one more if you wouldnt mind?

geerky42 (geerky42):

haha sure

OpenStudy (anonymous):

sure :D

OpenStudy (anonymous):

CooooooooOOOOOOL k

OpenStudy (anonymous):

What is the sum of the geometric sequence 1, -4, 16, ... if there are 7 terms? -5,461 5,461 -3,277 3,277

OpenStudy (anonymous):

\[a _{1} = 1\] \[a _{2} = -4\] \[a _{3} = 16\] \[r = -4/1 = -4 \] \[r = 16/-4 = -4 \] so \[r = -4 \]

OpenStudy (anonymous):

This is not the answer its is just the basic things your suppose to know of the geometric sequence to start.

OpenStudy (anonymous):

\[a _{1} = 1 * -4 = -4 \] \[a _{2} = -4 * -4 = 16\] \[a _{3} = 16 * -4 = -64\] \[a _{4} = -64 * -4 = 256\] \[a _{5} = 256 * -4 = -1024\] \[a _{6} = - 1024 * -4 = 4096\] \[a _{7} = 4096 * -4 = -16384\]

OpenStudy (anonymous):

now add them up.

OpenStudy (anonymous):

3,227 :)

OpenStudy (anonymous):

Thank you I really appreciate it!

OpenStudy (anonymous):

no problem :)